OpenStudy (anonymous):

what is the value of cos(tan^-1, 1)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

find \(\tan^{-1}1\) first

ganeshie8 (ganeshie8):

we should memorize that value :- tan(pi/4) = 1 so, tan^-1(1) = pi/4

ganeshie8 (ganeshie8):

so, \(\cos(\tan^{-1}1) = \cos(\pi/4) = ?\)

OpenStudy (anonymous):

is it 2?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

cos value lies between -1 and +1 so 2 is not an answer

ganeshie8 (ganeshie8):

look at that table, and see if u can find the cos(pi/4) value :)

OpenStudy (anonymous):

square root of 2?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

hmm no. here it is :- \(\large \cos(\pi/4) = \frac{\sqrt{2}}{2}\)

ganeshie8 (ganeshie8):

it helps in ur rest of course in trig, if u can memorize that table

ganeshie8 (ganeshie8):

just memorize the sin row, u can generate all other rows from sin